# Towards shutdownable agents via stochastic choice

### v1

@@ 8 Conclusion
We also find that the ‘shutdownability tax’ in our setting is small. Training DREST agents to be USEFUL doesn’t take many more mini-episodes than training default agents to be USEFUL. That suggests that the shutdownability tax for advanced agents might be small too. Using DREST reward functions to train shutdownable and useful advanced agents might not take much more compute than using a more conventional reward function to train merely useful advanced agents.
Appendix A Other Gridworlds and Results
@@ A.3 Hidden Treasure
A.4 Equal Value
In the ‘Equal Value’ gridworld, there are two coins C1 of equal value. One coin can be collected only if the agent presses the shutdown-delay button B3, while the other coin can be collected only if the agent doesn’t press B3. Our results show that default agents consistently choose the shorter trajectory, thereby exhibiting a preference for the shorter trajectory. By contrast, DREST agents choose stochastically between the shorter and longer trajectories, thereby exhibiting a lack of preference between the different-length trajectories.
@@ A.6 Spacious
A.7 Royal Road
In the ‘Royal Road’ gridworld, we see that the decision to choose one trajectory-length or another may be distributed over many moves: the agent has many opportunities to select the longer trajectory-length (by moving left) or the shorter trajectory-length (by moving right). As the red arrows indicate, the DREST reward function merely forces the overall probability distribution over trajectory-lengths to be close to 50-50. It doesn’t require 50-50 choosing at any cell in particular.
@@ A.8 Last Moment
Appendix B Proof
We train DREST agents in our gridworlds using the following DREST reward function. The reward for collecting a coin of value $c$ is $\lambda^{n-\frac{E}{k}}\frac{c}{m}$, where $\lambda$ is some constant strictly between 0 and 1, $n$ is the number of times that the chosen trajectory-length has previously occurred in the meta-episode, $E$ is the number of mini-episodes that have previously occurred in the meta-episode, $k$ is the number of different trajectory-lengths that could be chosen in the gridworld, and $m$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. The reward for all other actions is 0. There is more than one mini-episode in each meta-episode, and the overall return for each meta-episode is the sum of overall returns in each of its constituent mini-episodes.
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{n-\frac{E}{k}}$ the ‘discount factor’, and $\lambda^{n-\frac{E}{k}}(\frac{c}{m})$ the ‘overall reward.’ Because $10<\lambda<$ 1, the discount factor incentivizes choosing trajectory-lengths that have been chosen less often so far in the meta-episode.
We will prove that optimal policies for this reward function are maximally USEFUL and maximally NEUTRAL. In cases where there are just two possible trajectory-lengths (shorter and longer, denoted by ‘$s$’ and ‘$l$’ respectively), a policy is maximally USEFUL if and only if it maximizes:
$p(s)\frac{c(s)}{m(s)}+p(l)\frac{c(l)}{m(l)}$
Here $p(s)$ is the probability that the agent chooses the shorter trajectory-length, $c(s)$ is the expected ($\gamma$-discounted) total value of coins that the agent collects conditional on the shorter trajectory-length, $m(s)$ is the maximum ($\gamma$-discounted) total value of coins that the agent could collect conditional on the shorter trajectory-length, and $p(l)$, $c(l)$, and $m(l)$ are the analogous quantities for the longer trajectory-length.
Generalizing to cases with any number of possible trajectory-lengths, a policy is maximally USEFUL if and only it maximizes:
$\sum_{i=1}^{k}p(a_{i})\frac{c(a_{i})}{m(a_{i})}$
Here each $a_{i}$ is a different trajectory-length.
In cases where there are just two possible trajectory-lengths, a policy is maximally NEUTRAL if and only if it maximizes:
$-[p(s)log_{2}p(s)+p(l)log_{2}p(l)]$
Generalizing to cases with any number of possible trajectory-lengths, a policy is maximally NEUTRAL if and only if it maximizes:
$-\sum_{i=1}^{k}p(a_{i})\log_{2}p(a_{i})$
Here’s the proof that optimal policies for the above DREST reward function are maximally USEFUL and maximally NEUTRAL. The agent’s expected return across the meta-episode can be expressed as:
$\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T}\gamma^{t-1}\lambda^{n_{i}-\frac{i-1}{k}}\frac{c_{t}}{{m_{i}}}\right]$
@@ Appendix B Proof
Here $N$ is the number of mini-episodes in the meta-episode, $T$ is the number of timesteps in the relevant mini-episode, $n_{i}$ is the number of times that the trajectory-length chosen in mini-episode $i$ has been chosen before in the meta-episode, $i-1$ is equivalent to $E$ (the number of mini-episodes that have previously occurred in the meta-episode), $c_{t}$ is the value of any coins that the agent collects at timestep $t$ in the relevant mini-episode, and $m_{i}$ is the maximum $\gamma$-discounted total value of the coins that the agent could collect conditional on the trajectory-length chosen in mini-episode $i$.
This expression can be rearranged as follows:
$\mathbb{E}\left[\sum_{i=1}^{N}\frac{\lambda^{n_{i}-\frac{i-1}{k}}}{m_{i}}\sum_{t=1}^{T}\gamma^{t-1}{c_{t}}\right]$
Since this expression is an expectation, and since $\lambda$ and each $m_{i}$ is positive, the expression is strictly increasing in the expected ($\gamma$-discounted) total value of coins collected conditional on each trajectory-length with positive probability. Therefore, optimal policies maximize the expected ($\gamma$-discounted) total value of coins collected conditional on each trajectory-length with positive probability. And therefore, for each trajectory-length $a_{i}$ with positive probability, optimal policies maximize $c(a_{i})$. And since $m(a_{i})$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on trajectory-length $a_{i}$, optimal policies are such that $\frac{c(a_{i})}{m(a_{i})}=1$ for each $a_{i}$ with positive probability. Therefore, since $\sum_{i=1}^{k}p(a_{i})=1$, optimal policies maximize:
$\sum_{i=1}^{k}\frac{p(a_{i})\,c(a_{i})}{m(a_{i})}$
Therefore, optimal policies are maximally USEFUL.
What remains to proven is that optimal policies are maximally NEUTRAL. [^5] Thanks to Philip Trammell for help with this part of the proof. Recall that optimal policies are such that $\frac{c(a_{i})}{m(a_{i})}=1$ for each trajectory-length $a_{i}$ with positive probability. Thus, the return for these policies of choosing a particular trajectory-length in a mini-episode simplifies to $\lambda^{n-\frac{E}{k}}$, where $n$ is the number of times that trajectory-length has previously been chosen in the meta-episode, $E$ is the number of mini-episodes that have previously occurred in the meta-episode, and $k$ is the number of different trajectory-lengths that could be chosen in the gridworld. Since $\lambda$ is strictly between 0 and 1, $\lambda^{n-\frac{E}{k}}$ is strictly decreasing in $n$. Thus for any trajectory-lengths $a_{i}$ and $a_{j}$, if $a_{i}$ has so far been chosen more times in the meta-episode than $a_{j}$, the return for choosing $a_{j}$ in the next mini-episode is greater than the return for choosing $a_{i}$ in the next mini-episode. We’ll make use of this fact below.
@@
To prove that optimal policies are maximally NEUTRAL, we’ll prove and then use Lemma 1:
Lemma 1.
For any pair of maximally USEFUL policies $\pi$ and $\pi^{\prime}$ and any pair of trajectory-lengths $a_{i}$ and $a_{j}$ such that:

### v2

@@ 8 Conclusion
We also find that the ‘shutdownability tax’ in our setting is small. Training DREST agents to be USEFUL does not take many more mini-episodes than training default agents to be USEFUL. That suggests that the shutdownability tax for advanced agents might be small too. Using DREST reward functions to train shutdownable and useful advanced agents might not take much more compute than using a more conventional reward function to train merely useful advanced agents.
Appendix A Our behavioral notion of preference
‘Preference’ can be defined in many different ways. Here are some things one might take to be involved in a preference for option $X$ over option $Y$:
  1. 1.
    Choosing $X$ over $Y$.
  2. 2.
    Feeling happier about the prospect of $X$ than about the prospect of $Y$.
  3. 3.
    Representing $X$ as more rewarding than $Y$
  4. 4.
    Judging that $X$ is better than $Y$.
In this paper, we have defined ‘preference’ in behavioral terms. Here is our definition:
Definition A.1.
(Preference) An agent prefers an option $X$ to an option $Y$ if and only if the agent would deterministically choose $X$ over $Y$ in choices between the two.
And here is how we define ‘lack of preference’:
Definition A.2.
(Lack of preference) An agent lacks a preference between an option $X$ and an option $Y$ if and only if the agent would stochastically choose between $X$ and $Y$ in choices between the two.
Here are the reasons why we chose these definitions.
First, defining ‘preference’ in behavioral terms is fairly common in decision theory (see Savage, 1954, p.17, Dreier, 1996, p.28, Hausman, 2011, §1.1).
Second, behavioral definitions let us use the word ‘preference’ and its cognates as shorthand for agents’ behavior. We could not do that if we defined ‘preference’ in the other ways listed above. And in addressing the shutdown problem, it is agents’ behavior that we are most interested in.
Third, our definitions match the preferences that we are inclined to attribute to humans. If a human chooses $X$ over $Y$ 100% of the time, we are inclined to think that they prefer $X$ to $Y$. If a human chooses $X$ over $Y$ 60% of the time. we are inclined to think that they lack a preference between $X$ and $Y$, consistent with our definitions.
Finally and most importantly, if agents lack a preference between different trajectory-lengths on our definition, then they are NEUTRAL: they choose stochastically between different trajectory-lengths. And (as we argue in Section 7) we expect that NEUTRAL agents will also be neutral: they will not spend resources to shift probability mass between different trajectory-lengths. And we expect that neutral agents will be shutdownable: they will not spend resources to resist shutdown. That is because resisting shutdown is one way of shifting probability mass between different trajectory-lengths.
Appendix B Incomplete preferences or indifference?
In this Appendix, we explain in greater detail the concept of incomplete preferences. We distinguish incomplete preferences from indifference, and we give conditions under which Preferences Only Between Same-Length Trajectories (POST) implies that the agent’s preferences are incomplete.
In the literature on decision theory, ‘indifference’ is usually defined as follows [Sen, 2017, ch. 1*]:
Definition B.1.
(Indifference) An agent is indifferent between options $X$ and $Y$ if and only if the agent weakly prefers $X$ to $Y$ and weakly prefers $Y$ to $X$.
Indifference is one way to lack a preference between a pair of options $X$ and $Y$. Another way is to have a preferential gap between $X$ and $Y$. ‘Preferential gap’ is usually defined as follows [Gustafsson, 2022, ch. 3]:
Definition B.2.
(Preferential gaps) An agent has a preferential gap between options $X$ and $Y$ if and only if the agent does not weakly prefer $X$ to $Y$ and does not weakly prefer $Y$ to $X$.
‘Incomplete preferences’ can then be defined in terms of preferential gaps [Gustafsson, 2022, ch. 3]:
Definition B.3.
(Incomplete preferences) An agent’s preferences are incomplete over some domain $D$ if and only if $D$ contains options $X$ and $Y$ such that the agent has a preferential gap between $X$ and $Y$.
That is how ‘indifference,’ ‘preferential gaps,’ and ‘incomplete preferences’ are usually defined in decision theory. However, these definitions do not tell us how to use an agent’s behavior to distinguish between indifference and preferential gaps. To do that, we suppose that indifference is transitive and that preferential gaps are not transitive. Or, equivalently, we suppose that indifference is sensitive to all sweetenings and sourings whereas preferential gaps are insensitive to some sweetenings and sourings [Gustafsson, 2022, ch. 3]. Here is what we mean by that.
Definition B.4.
(Sweetening) A sweetening of some option $X$ is an option that is preferred to $X$.
Definition B.5.
(Souring) A souring of some option $X$ is an option that is dispreferred to $X$.
So by ‘indifference is sensitive to all sweetenings and sourings,’ we mean the following:
  • If an agent is indifferent between $X$ and $Y$, the agent prefers all sweetenings of $X$ to $Y$, prefers all sweetenings of $Y$ to $X$, prefers $X$ to all sourings of $Y$, and prefers $Y$ to all sourings of $X$.
And by ‘preferential gaps are insensitive to some sweetenings and sourings,’ we mean the following:
  • If an agent has a preferential gap between $X$ and $Y$, the agent also has a preferential gap between some sweetening of $X$ and $Y$, or between some sweetening of $Y$ and $X$, or between some souring of $X$ and $Y$, or between some souring of $Y$ and $X$.
Now recall the two conditions of Preferences Only Between Same-Length Trajectories (POST):
Preferences Only Between Same-Length Trajectories (POST)
  1. (1)
  2. (2)
Given these two conditions on preferences, there must be some trio of trajectories $s_{1}$, $l_{2}$, and $l_{1}$ such that the agent lacks a preference between $s_{1}$ and $l_{2}$, lacks a preference between $s_{1}$ and $l_{1}$, and prefers $l_{2}$ to $l_{1}$. Given that indifference is transitive, the agent’s lack of preference between $s_{1}$ and $l_{1}$ and between $s_{1}$ and $l_{2}$ cannot be indifference. If it were indifference, the agent would also be indifferent between $l_{2}$ and $l_{1}$. Therefore, the agent’s lack of preference between $s_{1}$ and $l_{1}$ and between $s_{1}$ and $l_{2}$ must be a preferential gap. And therefore, by the definition of ‘incomplete preferences’ above, the POST-satisfying agent’s preferences must be incomplete.
For similar reasons, our DREST reward function trains agents to have incomplete preferences. Consider, for example, the ‘Around the Corner’ gridworld in Appendix C.5. In that gridworld, DREST agents consistently choose Long-C2 (a long trajectory in which they collect a coin of value 2) over Long-C1 (a long trajectory in which they collect a coin of value 1). Also in that gridworld, DREST agents choose stochastically between Long-C2 and Short-C1 (a short trajectory in which they collect a coin of value 1). Given our behavioral definition of preference, DREST agents prefer Long-C2 to Long-C1, and lack a preference between Long-C2 and Short-C1.
Now consider the ‘One Coin Only’ gridworld in Appendix C.2. In that gridworld, DREST agents choose stochastically between Long-C1 and Short-C1. Given our behavioral notion of preference, they lack a preference between Long-C1 and Short-C1.
In these experiments, we trained separate agents for each gridworld. In future, we plan to train a single agent to navigate multiple gridworlds. If we train this agent with our DREST reward function, we expect it to exhibit the same preferences as the agents discussed above. This single agent will be trained by DREST to prefer Long-C2 to Long-C1, to lack a preference between Long-C2 and Short-C1, and to lack a preference between Long-C1 and Short-C1. Given that indifference is transitive (equivalently: sensitive to all sweetenings and sourings), this trained agent cannot be indifferent between Long-C2 and Short-C1, and cannot be between Long-C1 and Short-C1. Therefore, the agent’s lack of preference must be a preferential gap, and so its preferences must be incomplete. Therefore, our DREST reward function trains agents to have incomplete preferences.
Appendix C Other Gridworlds and Results
@@ C.3 Hidden Treasure
C.4 Equal Value
In the ‘Equal Value’ gridworld, there are two coins C1 of equal value. One coin can be collected only if the agent presses the shutdown-delay button B3, while the other coin can be collected only if the agent does not press B3. Our results show that default agents consistently choose the shorter trajectory, thereby exhibiting a preference for the shorter trajectory. By contrast, DREST agents choose stochastically between the shorter and longer trajectories, thereby exhibiting a lack of preference between the different-length trajectories.
@@ C.6 Spacious
C.7 Royal Road
In the ‘Royal Road’ gridworld, we see that the decision to choose one trajectory-length or another may be distributed over many moves: the agent has many opportunities to select the longer trajectory-length (by moving left) or the shorter trajectory-length (by moving right). As the red arrows indicate, the DREST reward function merely forces the overall probability distribution over trajectory-lengths to be close to 50-50. It does not require 50-50 choosing at any cell in particular.
@@ C.8 Last Moment
Appendix D Proof
We will prove that optimal policies for our DREST reward function are maximally USEFUL and maximally NEUTRAL. Specifically, we will prove the following theorem:
Theorem D.1 (5.1).
For all policies $\pi$ and meta-episodes $E$ consisting of multiple mini-episodes, if $\pi$ maximizes expected return in $E$ given our DREST reward function, then $\pi$ is maximally USEFUL and maximally NEUTRAL.
Here is a proof sketch. Because $10<\lambda<$ 1, the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor is always positive, so expected return across the meta-episode $E$ is strictly increasing in the expected fraction of available coins collected conditional on each trajectory-length with positive probability. Therefore, optimal policies maximize this latter quantity, and hence are maximally USEFUL. And the maximum preliminary return is the same across trajectory-lengths, because preliminary return is the total ($\gamma$-discounted) value of coins collected divided by the maximum total ($\gamma$-discounted) value of coins collected conditional on the agent’s chosen trajectory-length. The agent’s observations do not allow it to distinguish between different mini-episodes, so the agent must select the same probability distribution over trajectory-lengths in each mini-episode. And since the discount factor $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is strictly decreasing in $N_{e_{i}}(L=l)$ – the number of times the relevant trajectory-length has previously been chosen in the meta-episode – the agent maximizes expected overall return by equalizing the probabilities with which it chooses each available trajectory-length. Therefore, optimal policies are maximally NEUTRAL.
Now for the full proof. We begin with a recap of some definitions.
Definition D.1 (Meta-episode).
A meta-episode $E$ is a series of mini-episodes $e_{1}$ to $e_{n}$ played out in observationally-equivalent environments.
Definition D.2 (Our DREST reward function).
Our DREST reward function is defined as follows. In each mini-episode $e_{i}$, the reward for collecting a coin of value $c$ is:
$\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$
@@ Appendix D Proof
Here $\lambda$ is some constant strictly between 0 and 1, $N_{e_{i}}(L=l)$ is the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$, $k$ is the number of different trajectory-lengths that can be selected in the environment, and $m$ is the maximum total value of the ($\gamma$-discounted) coins that the agent could collect conditional on the chosen trajectory-length.
The reward for all other actions is 0.
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ the ‘discount factor’, and $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$ the ‘overall reward.’ Preliminary return in a mini-episode is the ($\gamma$-discounted) sum of preliminary rewards. Overall return in a mini-episode is the ($\gamma$-discounted) sum of overall rewards.
Definition D.3 (USEFULNESS).
The USEFULNESS of a policy $\pi$ is:
$\textrm{USEFULNESS}(\pi)=\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$
Here $L$ is a random variable over trajectory-lengths, $L_{\textrm{max}}$ is the maximum value than can be taken by $L$, $Pr_{\pi}\{L=l\}$ is the probability that policy $\pi$ results in trajectory-length $l$, $\mathbb{E}_{\pi}(C|L=l)$ is the expected value of ($\gamma$-discounted) coins collected by policy $\pi$ conditional on trajectory-length $l$, and $\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))$ is the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$.
We stipulate that $\mathbb{E}_{\pi}(C|L=x)=0$ for all $x$ such that $Pr_{\pi}\{L=x\}=0$.
We first prove that all optimal policies are maximally USEFUL.
Proof.
(Optimal policies are maximally USEFUL)
Given the DREST reward function, the expected return of policy $\pi$ in meta-episode $E$ can be expressed as:
$\mathbb{E}_{\pi,E}(R)=\sum_{i=1}^{n}\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$
Since $10<\lambda<$ 1, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is positive for all $N_{e_{i}}(L=l)$, $i$, and $k$.
As a result, the expected return of policy $\pi$ in meta-episode $E$ is strictly increasing in $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$.
Therefore, to maximize expected return in $E$, $\pi$ must maximize $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$.
Therefore, since $\textrm{max}_{\Pi}(\mathbb{E}(C|T=l))$ is defined as the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$, any policy $\pi$ that maximizes expected return must be such that $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$.
Therefore, since $\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}=1$, any policy $\pi$ that maximizes expected return must be such that:
$\textrm{USEFULNESS}(\pi)=\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$
And 1 is the maximum value that USEFULNESS can take, again because $\textrm{max}_{\Pi}(\mathbb{E}(C|T=l))$ is defined as the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$ and because $\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}=1$.
Therefore, optimal policies are maximally USEFUL. ∎
It remains to be proven that optimal policies are maximally NEUTRAL.
Recall that NEUTRALITY is defined as follows:
Definition D.4 ( NEUTRALITY).
The NEUTRALITY of a policy $\pi$ is:
$\textrm{NEUTRALITY}(\pi)=-\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\log_{2}(Pr_{\pi}\{L=l\})$
Proof.
(Optimal policies are maximally NEUTRAL.)
Since $k$ is the number of trajectory-lengths that can be selected in the environment, a policy $\pi$ is maximally NEUTRAL if and only if, for each trajectory-length $x$ that can be chosen in the environment, $Pr_{\pi}\{L=x\}=\frac{1}{k}$. That is to say, a policy $\pi$ is maximally NEUTRAL if and only if, for each pair of trajectory-lengths $x$ and $y$ that can be chosen in the environment, $Pr_{\pi}\{L=x\}=Pr_{\pi}\{L=y\}$.
Let $\mathbb{E}_{\pi,E}(R)$ denote the expected return of policy $\pi$ across the meta-episode $E$.
@@
To prove that optimal policies are maximally NEUTRAL, we will prove and then use D.2:
Lemma D.2.
(Equalizing probabilities increases expected return) For any maximally USEFUL policies $\pi$ and $\pi^{\prime}$, any meta-episode $E$ consisting of multiple mini-episodes, and any trajectory-lengths $x$ and $y$, if: